# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

### v1

@@ 5.4 Ablation Studies
Effects of RCORE components.
In Table 4 (a), we investigate the effect of each component in our RCORE framework. When used individually, TORC loss is the most effective by providing explicit temporal modeling objectives. Notably, it significantly improves verb recognition performances across both seen and unseen compositions (63.6% vs. 65.7% on seen compositions and 54.4% vs. 56.8% on unseen compositions). Importantly, we can achieve the highest ZS-CAR performance by employing all three components together during training. This synergistic improvement is achieved by effectively tackling dataset sparsity with VOCAMix, reinforcing temporal order modeling with TORC, and alleviating co-occurrence bias with Margin loss.
Effects of TORC components.
@@
Table 4 (b) evaluates the roles of $L_{\cos}$, which enforces discrimination of opposite temporal semantics, and $L_{\text{ent}}$, which prevents reliance on static cues. We show the each component of TORC can encourage performance improvement. Each term individually improves performance, while using both together yields the largest gains in unseen verb accuracy.
Effects of VOCAMix.
In Table 4 (c), we validate the effectiveness of our proposed composition-aware VOCAMix by comparing it with widely adopted mixing augmentations. We argue that conventional image-based augmentations are insufficient for composition verb-object learning. Specifically, CutMix [38] allows for mixing backgrounds with corresponding labels and Mixup [41] blends verbs and objects concurrently, failing to preserve the temporal difference information that is essential for verb recognition. As a result, these methods yield a lower H.M. score than the baseline. In contrast, our VOCAMix, aiming to generate realistic unseen composition samples, shows modest performance gains over the baseline. Furthermore, we observe that mixing a static frame into the estimated foreground region shows additional performance boost (+0.2 points in the H.M.) over applying it to the entire background (denoted as ‘VOCAMix w/o F.E.’).
Effects of Margin loss.
In Table 4 (d), we validate the effectiveness of our proposed Margin loss. To mitigate the model’s reliance on dataset co-occurrence priors while maintaining baseline performance, we employ our method and validate its effectiveness by comparing it with the widely-used large-margin softmax loss. The large-margin softmax (L-softmax) [19] loss is implemented by multiplying the angle $\theta$ between the input feature vector and the target classifier by a margin $m$ (e.g., $\cos(m\theta)$), thereby forcing the model to learn stricter decision boundaries with narrower angular margins. We apply this L-softmax loss into composition classification loss, consistent with the implementation schedule of our Margin loss. Compared to the L-softmax [19] approach, our proposed Margin loss successfully reduces FCP ratio while preserving performance. This demonstrates the effectiveness of our strategy, which selectively targets frequent-but-incorrect logits not the entire logits.
@@ 6 Conclusions
We revisit zero-shot compositional action recognition and show that existing methods fail mainly due to object-driven shortcuts, stemming from compositional sparsity and the asymmetric difficulty of learning verbs versus objects. We also find that the widely used closed-world, test-tuned evaluation protocol conceals this failure, overstating model generalization. We introduce RCORE, a simple framework that targets these issues directly: VOCAMix enlarges the effective compositional support, and TORC loss enforces temporal order sensitivity to reduce reliance on static object cues. Across two benchmarks and under a realistic open-world, unbiased evaluation setting, RCORE consistently improves unseen composition recognition performances, indicating genuine compositional reasoning. Our results highlight that both proper evaluation and shortcut-aware training are essential for progress in compositional video understanding.
@@ Acknowledgments
In this appendix, we provide comprehensive implementation/dataset/method details and quantitative/qualitative results to complement the main paper. We organize the appendix as follows:
@@
    Details on EK100-com dataset (Section A).
@@
    Complete implementation details (Section B).
@@
    Additional evidence of object-driven shortcuts (Section C).
@@
    Additional results (Section D).
Appendix A Details on EK100-com dataset
@@
In this section, we provide details about our curated ZS-CAR benchmark, EPIC-KITCHENS-100-composition (EK100-com). We construct EK100-com by repurposing EPIC-KITCHENS-100 (EK100) [7] following the same protocol of constructing Sth-com [16]. In particular, we use the original training (67217 samples) and validation (9668 samples) split of EK100 [7] and split them as follows: (1) we filter the data to include only compositions with more than five samples each. (2) we ensure that all compositions present in the validation set also exist in the training set. After initial filtering, the training set comprises 62,790 samples with 1,331 unique compositions and the validation set contains 8,657 samples, which consist of 833 seen compositions and 3 unseen compositions. (3) we randomly select half of the compositions from both the training and validation sets and swap these subsets. (4) Finally, we split the validation and test set with a ratio of $3:4$. We provide the statistics of EK100-com in Table 5 and Figure 7, and some seen/unseen composition examples of EK100-com in Figure 8.
@@
Beyond expanding the pool of publicly available ZS-CAR benchmarks, our repurposed EK100-com exhibits two key properties that make it a compelling ZS-CAR benchmark. First, unlike Something-Something V2 [12], where videos may involve multiple objects associated with a verb, videos in EK100 [7] depict actions consisting of a single verb and a single object. Therefore, the training data of EK100-com provides explicit object supervision for the models. Second, since EK100 [7] has long-tailed distribution, the label coverage ratio of EK100-com is 7.5%, which is much lower than that of Sth-com [16] (12.8%). This makes EK100-com a challenging benchmark, suitable for evaluating whether a model can overcome the risk of shortcut learning induced by data sparsity.
(a) Top-30 frequent compositions
(b) Bottom-30 frequent compositions
@@
(a) Seen composition examples
(b) Unseen composition examples
@@
Appendix B Complete Implementation Details
B.1 Implementation details on our diagnosis
@@
In this section, we provide comprehensive details of experimental setup and implementation details of our diagnosis for Zero-Shot Compositional Action Recognition (ZS-CAR) in Section 3 in the main paper. All notations used in this supplementary document are consistent with those defined in the main paper. As mentioned in the main paper, we adopt the state-of-the-arts model, C2C [16] which uses a conditional learning approach, as the existing ZS-CAR baseline for diagnosis.
Details about our diagnosis in Figure 2 (a).
@@
In Figure 2 (a) in the main paper, we empirically demonstrate the asymmetric learning difficulty between verbs and objects. We train a randomly initialized ViT [10] with joint spatio-temporal attention on our selected subset of the Sth-com [16] dataset. We aim to select a subset that maximizes the number of samples while maintaining a sufficient number of distinct labels. As a result, the selected subset is composed of 10 verb classes: ‘Moving sth and sth away from each other’, ‘Moving sth and sth closer to each other’, ‘Moving sth away from sth’, ‘Moving sth closer to sth’, ‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, ‘Pushing sth from right to left’, ‘Pushing sth so that it falls off the table’, and ‘Pushing sth so that it slightly moves’; and 10 object classes: ‘Ball’, ‘Battery’, ‘Book’, ‘Bottle’, ‘Coin’, ‘Glass’, ‘Knife’, ‘Paper’, ‘Remote’, and ‘Spoon’. We construct the training set consisting of 2,555 samples from the original Sth-Com [16] training and validation samples, while the validation set comprises 260 samples from the original Sth-Com [16] test set. For training, we employ a multi-task learning approach using separate linear classifiers for verbs and objects, taking the ViT’s CLS token as input. In the main paper, we show that objects are easier to learn than verbs by monitoring the learning curves of a randomly initialized ViT. Furthermore, we substantiate in Section C demonstrating that this asymmetric learning difficulty leads the model to exploit object-driven shortcuts in compositional video understanding.
Details about our diagnosis in Figure 2 (b).
@@
To precisely control the sample counts for each verb-object pair, we select a $4\times 4$ toy set from the Sth-Com [16] dataset. This set contains four verb labels (‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, and ‘Pushing sth from right to left’) and four object labels (Book, Remote, Glass, Bottle). The training set consists of the four diagonal pairs (refer to Figure 2 (b)) with 40 samples per pair. The test set comprises two distinct splits: seen compositions (25 samples per pair) and unseen compositions (30 samples per pair). We use 8 frames for input videos.
Details about our diagnosis in Figure 3.
@@
In Figure 3 in the main paper, we present the learning curves of the baseline model [16] trained on the Sth-com [16]. We further present diagnostic results for both the baseline and RCORE on EK100-com in Figure 9. We define two diagnostic metrics, False Seen Prediction (FSP) ratio and False Co-occurrence Prediction (FCP) ratio. FSP measures the proportion of unseen validation samples incorrectly classified as seen compositions:
@@
$\textit{False Seen Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}})},$ (7)
@@
where $\mathbb{I}(\cdot)$ is the indicator function. FCP further quantifies the fraction of these false-seen predictions that correspond to frequently co-occurring verb–object pairs in the training data:
@@
$\textit{False Co-oc. Prediction}=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{freq}})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}})},$ (8)
where $\mathbb{Y}_{\text{freq}}$ denotes the set of frequently observed verb–object pairs in the training set. $\mathbb{Y}_{\text{freq}}$ is defined with the frequency distribution of the training set $P_{\mathbb{D}_{train}}$ as follows:
$\mathbb{Y}_{\text{freq}}$ $=\mathbb{Y}_{\text{freq}}^{O\mid V}\cap\mathbb{Y}_{\text{freq}}^{V\mid O},$ (9)
where $\mathbb{Y}_{\text{freq}}^{O\mid V}$ is defined as the set of pairs where the conditional probability of an object given a verb in the training data is greater than $\mu+\sigma$. Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the conditional probabilities calculated on the training set. The $\mathbb{Y}_{\text{freq}}^{V\mid O}$ is defined similarly for the conditional probability of a verb given an object. $(\mu_{O|V}+\sigma_{O|V},\mu_{V|O}+\sigma_{V|O})$ are calculated as $(0.13,0.19)$ for the Sth-com [16] dataset and $(0.23,0.40)$ for EK100-com.
@@
Figure 9: Learning curve of the baseline and RCORE on the EK100-com dataset. RCORE suppresses the increase of the FCP ratio during training, effectively narrowing the performance gap between seen and unseen composition validation accuracies.
B.2 Implementation details on RCORE
Experimental Setup.
We implement our method using PyTorch and conduct all experiments on 8 NVIDIA Tesla V100 GPUs. For video preprocessing, we uniformly sample frames from each video to generate clips of size $3\times 16\times 224\times 224$ (channels $\times$ frames $\times$ height $\times$ width).
Model Architecture.
We utilize CLIP-B/16 [28] as our visual/text backbone. Following prior work [16], we adopt AIM [37], a parameter-efficient spatio-temporal adapter for CLIP [28], as the video encoder. We also construct the verb encoder with two temporal convolution layers with ReLU activations, while the object encoder employs temporal average pooling followed by a two-layer MLP with ReLU activations. We use CoOp-style [43] learnable text prompts for all models.
Training Strategy.
We train all models, including baselines and RCORE, for 30 epochs with a total batch size of 128. The base learning rate is set to $5\times 10^{-4}$ for Sth-com [16] and $1\times 10^{-4}$ for EK100-com, initialized with a 3-epoch warmup. Following prior work [16], we adopt closed-world training setting. We empirically find that setting the softmax temperature to 0.07, same as the implementation of CLIP [28], results in better ZS-CAR performance than the value of 0.01, used in prior work [16]. Therefore, we adopt this temperature value for both C2C [16] and RCORE, except for the case of AIM [37] with independent modeling. For VOCAMix, we employ the unsupervised foreground estimation algorithm FAME [9] to identify high-motion regions. FAME [9] isolates the moving foreground from background regions based on frame differences and color statistics. To generate the mixed samples, we first sample the mixing coefficient from a $\text{Beta}(2.0,2.0)$ distribution, following prior work [38, 41]. We then scale it by 0.2 for Sth-com [16] and 0.5 for EK100-com to obtain the final blending coefficient $\lambda$. We summarize the hyperparameters used in this work in Table 6.
Loss configuration.
For the total loss, we set the weights for the verb and object components to $\alpha=0.2$ and for the composition loss to $\beta=1.0$, following prior work [16]. As demonstrated in Table 9 (g) and (h), we empirically find it beneficial to linearly warm up the weights for the auxiliary losses: $\gamma$ is increased from 0.0 to 1.0 over 5 epoch to 10 epoch, and $\delta$ from 0.0 to 0.8 over epochs 15 epoch to 20 epoch.
Inference.
During inference, we adopt a single-view, single-crop protocol. For bias calibration, we select the bias term that maximizes the harmonic mean of the accuracies for seen and unseen compositions.

### v2

@@ 5.4 Ablation Studies
Effects of RCORE components. In Table 5 (a), using TORC alone yields the largest gain in verb generalization, improving verb@unseen-comp by 3.5 points, and gain in unseen composition accuracy by 3.8 points, compared to the baseline without CPR and TORC (C2C [li2024c2c]). Using CPR alone primarily improves unseen composition accuracy by 3.1 points, while reducing seen composition accuracy 4.2 points, indicating a seen–unseen trade-off. Combining CPR and TORC achieves the best overall results, improving unseen composition accuracy by +5.0 points and composition H.M. by +2.7 points, suggesting complementary benefits of mitigating co-occurrence bias and enforcing temporal sensitivity.
@@
Effects of TORC components. Table 5 (b) evaluates the roles of $L_{\cos}$, which enforces discrimination of opposite temporal semantics, and $L_{\text{ent}}$, which prevents reliance on static cues. Each term individually improves performance, while using both terms yields the largest gains in the verb@unseen-comp (+3.5 points) and unseen composition accuracy (+3.8 points).
Table 5: Ablation study on the Sth-com [li2024c2c] validation set. We use C2C [li2024c2c] with the CLIP [radford2021clip] backbone as our baseline, using the unbiased open-world setting. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.). V@S and V@U denotes verb@seen-comp and verb@unseen-comp, while O@S and O@U denotes object@seen-comp and object@unseen-comp accuracies.
(a) Effects of RCORE components.
CPR TORC Verb Object Composition
V@S V@U O@S O@U Seen Unseen H.M.
65.26 54.02 67.21 56.81 47.31 30.14 36.83
63.95 54.05 63.97 59.52 43.09 33.22 37.51
64.93 57.49 66.20 57.23 46.28 33.92 39.15
67.06 58.13 63.67 58.66 45.00 35.16 39.48
(b) Effects of TORC components. $L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) V@S V@U O@S O@U Seen Unseen H.M. 65.26 54.02 67.21 56.81 47.31 30.14 36.83 64.15 55.96 67.71 55.97 47.13 31.11 37.48 64.84 54.82 67.57 56.23 47.46 30.84 37.39 64.93 57.49 66.20 57.23 46.28 33.92 39.15
(c) Effects of label space in CPR. Label Space Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M. Closed-world ($\mathbb{Y}_{\text{seen}}$) 64.61 54.26 67.21 57.46 46.78 30.80 37.14 Full Open-world ($\mathbb{Y}^{V}\times\mathbb{Y}^{O}$) 67.36 45.01 68.49 52.29 50.72 17.34 25.85 Batch-adaptive ($\mathbb{Y}_{\text{exp}}$) 64.17 54.55 63.60 58.70 42.74 33.05 37.28
(d) Effects of penalty set in CPR. Penalty Set $\mathbb{Y}_{\text{pen}}$ Choice Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M. None (w/o $L_{\text{CPR}}$) 64.17 54.55 63.60 58.70 42.74 33.05 37.28 All Hard Negatives ($\mathbb{Y}_{\text{hard}}$) 61.95 57.22 47.99 49.05 31.85 30.82 31.33 Frequent Hard Negatives ($\mathbb{Y}_{\text{hard}}\cap\mathbb{Y}_{\text{freq}}$) 65.36 55.86 63.20 57.62 43.60 33.16 37.67
Effects of label space in CPR. In Table 5 (c), we validate the batch-adaptive label space expansion in CPR. Compared to closed-world training over $\mathbb{Y}_{\text{seen}}$, our batch-adaptive label space improves unseen composition accuracy (+2.3 points) and H.M. by injecting supervision for synthesized compositions while avoiding unstable optimization. In contrast, naively optimizing cross-entropy over the full $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ space severely hurts unseen composition accuracy by treating most unseen compositions as negatives throughout training.
Effects of penalty set in CPR. In Table 5 (d), we study the effect of the penalty set $\mathbb{Y}_{\text{pen}}$. Compared to disabling $L_{\text{CPR}}$, penalizing frequent hard negatives ($\mathbb{Y}_{\text{hard}}\cap\mathbb{Y}_{\text{freq}}$) yields the best trade-off, improving unseen composition accuracy and achieving the highest H.M. of 37.7%. In contrast, penalizing all hard negatives degrades object prediction and hurts overall composition performance, suggesting that restricting penalties to frequent confounders is critical.
@@ 6 Conclusions
In this work, we revisit ZS-CAR and identify object-driven shortcuts as a key failure mode, driven by compositional sparsity and asymmetric verb–object learning difficulty. Using our diagnostic metrics, we show that existing models overfit to co-occurrence priors and consequently generalize poorly to unseen compositions. We propose RCORE with CPR (synthesized supervision with frequent hard-negative regularization) and TORC (temporal-order sensitivity), and demonstrate consistent gains on two benchmarks under open-world evaluation without test-tuned calibration.
@@ Acknowledgments
In this appendix, we provide comprehensive evaluation-protocol/ dataset/implementation details and additional results to complement the main paper. We organize the appendix as follows:
@@
    Overview of our evaluation setting (Section 0.A).
@@
    Details on EK100-com dataset (Section 0.B).
@@
    Complete implementation details (Section 0.C).
@@
    Additional evidence of object-driven shortcuts (Section 0.D).
  5. 5.
    Additional results (Section 0.E).
Appendix 0.A Overview of our evaluation setting
In Figure 7 and Figure 8, we briefly describe our evaluation setting. In Zero-Shot Compositional Action Recognition (ZS-CAR), the compositional label space spans the Cartesian product of the verb and object primitives ($|\mathbb{Y}^{V}|\times|\mathbb{Y}^{O}|$). At test time, samples are categorized into seen and unseen compositions depending on whether they were observed during training, as shown in Figure 7 (a). Under the closed-world evaluation setting, the evaluation is restricted to compositions present within the dataset (i.e., the union of the training and test sets). Consequently, the model computes compositional logits exclusively for these predefined closed-world pairs, as shown in Figure 7 (b).
However, this closed-world setting suffers from the impractical assumption of requiring prior knowledge of the test set’s ground-truth labels. Therefore, we adopt the open-world evaluation setting as our primary protocol. By evaluating across the entire compositional space and computing logits for all possible combinations, as shown in Figure 7 (b), the open-world setting enables the evaluation of various possible-but-unseen compositions encountered in real-world applications, where the test distribution is unknown.
(a) Closed-world and open-world evaluation settings
(b) An example of closed/open-world evaluation
Figure 7: Closed-world and open-world evaluation protocols in ZS-CAR. (a) (Left) Primitives are learned from a subset of training compositions (Train Pairs). (Middle) Closed-world evaluation restricts inference to dataset-present pairs (Train Pairs and Test Seen/Unseen Pairs), explicitly ignoring others as Masked Pairs. (Right) Open-world evaluation spans the entire compositional space, requiring the model to navigate all Possible Unseen Pairs without artificial masking. (b) In the closed-world setting, the Possible Unseen Pair—not appearing in the dataset—(Close, Window) becomes a Masked Pair, and its logit is ignored. Consequently, the next valid class (Open, Window) is selected as the top-1 prediction. This closed-world setting constrains the evaluation of numerous possible-but-unseen compositions. However, in the open-world setting, the model treats all pairs as valid, thus predicting (Close, Window) as the top-1 prediction.
As shown in Figure 8, we measure verb, object, and composition accuracies across seen and unseen splits to evaluate not only the model’s compositional recognition capability, but also how primitive recognition degrades on unseen compositions. Specifically, based on the model’s composition predictions, we verify whether the composition, verb, or object is predicted correctly, yielding three sets of metrics: (1) seen/unseen composition accuracy, (2) verb@seen/unseen-comp, and (3) object@seen/unseen-comp.
Figure 8: Evaluation metrics used in this work. We evaluate accuracies for verbs, objects, and compositions on seen and unseen composition splits separately. Based on the model’s composition predictions, we count the correct matches at each level to report (1) seen/unseen composition accuracy, (2) verb@seen/unseen-comp, and (3) object@seen/unseen-comp. For instance, predicting an unseen composition (Lift up, Spoon) as (Lift up, Pen) updates only the correct count for verb@unseen-comp (orange arrow).
Appendix 0.B Details on EK100-com dataset
@@
In this section, we provide details about our curated ZS-CAR benchmark, EPIC-KITCHENS-100-composition (EK100-com). We construct EK100-com by repurposing EPIC-KITCHENS-100 (EK100) [damen2022ek100] following the same protocol of constructing Sth-com [li2024c2c]. In particular, we use the original training (67,217 samples) and validation (9,668 samples) split of EK100 [damen2022ek100] and split them as follows: (1) we filter the data to include only compositions with more than five samples each. (2) we ensure that all compositions present in the validation set also exist in the training set. After initial filtering, the training set comprises 62,790 samples with 1,331 unique compositions and the validation set contains 8,657 samples, which consist of 833 seen compositions and 3 unseen compositions. (3) we randomly select half of the compositions from both the training and validation sets and swap these subsets. (4) Finally, we split the validation and test set with a ratio of $3:4$. We provide the statistics of EK100-com in Table 6 and Figure 9, and some seen/unseen composition examples of EK100-com in Figure 10.
@@
Beyond expanding the pool of publicly available ZS-CAR benchmarks, our repurposed EK100-com exhibits two key properties that make it a compelling ZS-CAR benchmark. First, unlike Something-Something V2 [goyal2017something], where videos may involve multiple objects associated with a verb, videos in EK100 [damen2022ek100] depict actions consisting of a single verb and a single object. Therefore, the training data of EK100-com provides explicit object supervision for the models. Second, since EK100 [damen2022ek100] has long-tailed distribution, the label coverage ratio—defined as the number of unique verb–object pairs present in the dataset relative to all possible combinations—–of EK100-com is 7.5%, which is much lower than that of Sth-com [li2024c2c] (12.8%). This makes EK100-com a challenging benchmark, suitable for evaluating whether a model can overcome the risk of shortcut learning induced by data sparsity.
@@
Top-30 frequent compositions
Bottom-30 frequent compositions
Seen composition examples
Unseen composition examples
@@
Appendix 0.C Complete Implementation Details
0.C.1 Implementation details on our diagnosis
@@
In this section, we provide comprehensive details of experimental setup and implementation details of our diagnosis for ZS-CAR in Section 3. All notations used in this appendix are consistent with those defined in the main paper. As mentioned in Section 3, we adopt the state-of-the-art model, C2C [li2024c2c], as the existing ZS-CAR baseline for diagnosis.
@@
Details about our diagnosis in Figure 2 (a). In Figure 2 (a) in the main paper, we empirically demonstrate the asymmetric learning difficulty between verbs and objects. We train a randomly initialized ViT [dosovitskiy202vit] with joint Spatio-Temporal (ST) attention on our selected subset of the Sth-com [li2024c2c] dataset. We aim to select a dense subset that eliminates sparsity by ensuring samples exist for every pair, maximizing the number of samples while maintaining a sufficient number of distinct labels. As a result, the selected subset is composed of 10 verb classes: ‘Moving sth and sth away from each other’, ‘Moving sth and sth closer to each other’, ‘Moving sth away from sth’, ‘Moving sth closer to sth’, ‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, ‘Pushing sth from right to left’, ‘Pushing sth so that it falls off the table’, and ‘Pushing sth so that it slightly moves’; and 10 object classes: ‘Ball’, ‘Battery’, ‘Book’, ‘Bottle’, ‘Coin’, ‘Glass’, ‘Knife’, ‘Paper’, ‘Remote’, and ‘Spoon’. We construct the training set consisting of 2,555 samples from the original Sth-Com [li2024c2c] training and validation samples, while the validation set comprises 260 samples from the original Sth-Com [li2024c2c] test set. For training, we employ a multi-task learning approach using separate linear classifiers for verbs and objects, taking the ViT’s CLS token as input.
Figure 11: Object-driven shortcuts with the random initialized model. Even with a randomly initialized model, a ViT-B [dosovitskiy202vit] with joint ST attention, object-driven shortcuts emerge: the model achieves high object accuracy, but verb accuracy on bias-conflicting unseen compositions drops below chance, indicating shortcut-driven verb failures.
@@
Details about our diagnosis in Figure 2 (b). To precisely control the sample counts for each verb-object pair, we select a $4\times 4$ toy set from the Sth-Com [li2024c2c] dataset. This set contains four verb labels (‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, and ‘Pushing sth from right to left’) and four object labels (Book, Remote, Glass, Bottle). As shown in Figure 2 and Figure 11, the training set consists of the four diagonal pairs with 40 samples per pair. The test set comprises two distinct splits: seen compositions (25 samples per pair) and unseen compositions (30 samples per pair). We use 8 frames for input videos. In Figure 11, we demonstrate that the object-driven verb shortcut learning observed with CLIP [radford2021clip] in the main paper emerges identically even in a randomly initialized model. This indicates that this phenomenon is not a side effect of pre-trained knowledge, but rather a structural issue inherent to ZS-CAR.
@@
Details about our diagnosis in Figure 3. In Figure 3, we present the learning curves of the baseline model [li2024c2c] trained on the Sth-com [li2024c2c]. We further present diagnostic results for both the baseline and RCORE on EK100-com in Figure 12. We define two misclassification ratios, False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP). FSP measures the proportion of unseen composition validation samples incorrectly classified as seen compositions:
@@
$\vskip-30.00005pt\textit{False Seen Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{seen}})},\vskip-5.0pt$ (9)
@@
where $\mathbb{I}(\cdot)$ is the indicator function. FCP further quantifies the fraction of these false-seen predictions that correspond to frequently co-occurring verb–object pairs $\mathbb{Y}_{\text{freq}}$ in the training data:
@@
$\vskip-10.00002pt\textit{False Co-oc. Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})}.\vskip-10.00002pt$ (10)
We further break down these misclassification ratios into the three cases: (i) Verb-collapse, where the predicted object is correct but the verb is incorrect; (ii) Object-collapse, where the predicted verb is correct but the object is incorrect; and (iii) Dual-collapse, where both verb and object are incorrect. We define them in FSP as follows:
Verb-collapse $=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})\cdot\mathbb{I}(\hat{\mathbf{y}}^{V}_{i}\neq\mathbf{y}^{V}_{i}\land\hat{\mathbf{y}}^{O}_{i}=\mathbf{y}^{O}_{i})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})}$ (11)
Object-collapse $=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})\cdot\mathbb{I}(\hat{\mathbf{y}}^{V}_{i}=\mathbf{y}^{V}_{i}\land\hat{\mathbf{y}}^{O}_{i}\neq\mathbf{y}^{O}_{i})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})}$ (12)
Dual-collapse $=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})\cdot\mathbb{I}(\hat{\mathbf{y}}^{V}_{i}\neq\mathbf{y}^{V}_{i}\land\hat{\mathbf{y}}^{O}_{i}\neq\mathbf{y}^{O}_{i})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})},$ (13)
where $\mathbf{y}^{C}_{i}=(\mathbf{y}^{V}_{i},\mathbf{y}^{O}_{i})$. We apply the identical formulation for FCP as well.
(a) FSP/FCP
(b) Composition accuracy
@@
Figure 12: Learning curve of the baseline and RCORE on EK100-com. RCORE (a) suppresses the increase of FCP during training, (b) effectively narrowing the performance gap between seen and unseen composition validation accuracies, compared to the baseline (C2C [li2024c2c]).
0.C.2 Implementation details on RCORE